OpenStudy (anonymous):

HELP ME PLEASE!!!! graph each function and its inverse. after describe the domain and range of each. f(x)=(x+3)^2

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

do you know how to graph \(\bf x^2\quad ?\)

OpenStudy (anonymous):

well i have a graphing calculator so it does it for me. i dont know how to find the inverse and the domain and range. i know when you graph x^2 you get a u shape graph

OpenStudy (anonymous):

Inverse of f(x) \[ g(x)=f^{-1}(x) =\sqrt x-3 \]

OpenStudy (anonymous):

so the inverse is \[\sqrt{x}-3\]

OpenStudy (jdoe0001):

\(\begin{array}{llll} f(x)=(x&+3)^2\\ &\uparrow \\ &\textit{horizontal shift of }x^2\textit{ graph to the left by 3 units} \end{array}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so how do i find the domain and range?

OpenStudy (anonymous):

The domain of is all real numbers. The domain of g is all real \( x \ge 0\)

OpenStudy (jdoe0001):

to find the inverse "relationship" of an expression, you'd just swap the variables about \(\bf f(x)=\color{red}{y}=(x+3)^2\qquad inverse\implies x=(\color{red}{y}+3)^2\\ \quad \\ \sqrt{x}=y+3\implies \sqrt{x}-3=y\) and solve for "y"

OpenStudy (anonymous):

ok i dont understand the domain and range still

OpenStudy (anonymous):

The domain is all x so that the function evaluated at x makes sense

OpenStudy (anonymous):

The range of a function h is all the values of h(x)

OpenStudy (jdoe0001):

the domain of a function is, all values that "x" can take, that is, the INPUT meaning, any values for "x" that produce a valid "y", not an undefined or extraneous values so for \(\bf f(x)=y=(x+3)^2\) "x" can pretty much take anything, thus \(\bf x \in \mathbb{R}\)

OpenStudy (anonymous):

@jdoe0001 can help you finish this problem

OpenStudy (jdoe0001):

for \(\bf \sqrt{x}-3=y\) the radical CANNOT be negative, because when it's negative you'd end up with an imaginary value, thus not a valid one so for \(\bf \sqrt{x}-3=y\) "x" cannot be negative, that is, has to always be either positive or 0 at the very least

OpenStudy (anonymous):

ok that will work but you lost me on the last part im so confused on the values of x and them being real numbers and all that. what is xr thing

OpenStudy (anonymous):

so the domain is 0 and up for the inverse?

OpenStudy (jdoe0001):

\(\begin{array}{llll} \mathbb{R} \implies \textit{all real numbers set}\\ \quad \\ \bf x\in \mathbb{R}\implies \textit{"x" belongs to all real numbers set} \end{array}\)

OpenStudy (jdoe0001):

0 and up, yes, I assume you've covered "imaginary numbers", so when you have a negative radicand, that's what you've got

OpenStudy (anonymous):

yes. so after that how do you find the range of the two functions

OpenStudy (jdoe0001):

..... well... notice your graph in the calculator for \(\bf f(x)=(x+3)^2\) what's its vertex?

OpenStudy (anonymous):

the vertex is -3??? i think

OpenStudy (jdoe0001):

-3? x= -3? y = -3?

OpenStudy (jdoe0001):

what's the (x,y) pair for the vertex?

OpenStudy (anonymous):

oh i forgot to put it in that form it is (-3,0)

OpenStudy (jdoe0001):

does the graph go below (-3, 0)? no, since that's the vertex, meaning that "y" values don't go pass 0, they do go ABOVE 0, not BELOW 0 so that's the range range= values for "y"

OpenStudy (jdoe0001):

you can also try graphing maybe \(\bf y=\sqrt{x}-3\) and you'll also see the range as well, by just looking at the graph

OpenStudy (anonymous):

so both x and y are values greater than or equal to 0 correct

OpenStudy (anonymous):

is that the domain and range for both functions?

OpenStudy (jdoe0001):

well.... for \(\bf f(x)=y=(x+3)^2\) the values for "y" don't go pass the vertex of (-3, 0) so the range goes ABOVE 0, not BELOW 0, that'd be the range for that one recall that for "x" or the domain was \(\bf x \in \mathbb{R}\)

OpenStudy (jdoe0001):

"so the range goes ABOVE 0, not BELOW 0" <---- so, what does that make the range?

OpenStudy (jdoe0001):